the tens digit of a 2 digit no. is 2 times its ones digit. if the digits are reversed,the sum of the original no. and the no. so formed is 66.find the number. .........................................help me plz i am stuck thank you.
Answer by oscargut(2103)

Letthe number = "ab"
a = 2b
10a+b + 10b+a = 66
11(a+b)= 66
a+b = 6
Then:
2b+b=6
3b=6
b = 2
a = 4
Answer: the number is 42
If you have any doubt or if you have more problems my e-mail is:
mthman@gmail.com